# Direct Sampling from Conditional Distributions by Sequential Maximum Likelihood Estimations

### v2

@@ 2.4 The approximate algorithm without using connection matrices
In this subsection, we propose to use Algorithm 1 with Steps 2 and 3 replaced by Steps 2’ and 3’, respectively. We call the resulting algorithm as the approximate direct sampling algorithm. From Lemma 2.4, when the MLE is not the UMVUE, i.e., it is not the rational MLE, then there exists a bias that disappears asymptotically. Thus the resulting algorithm is approximate in the sense that it is only asymptotically exact, not exact.
@@
For a log-linear model, the approximate algorithm obtained from Algorithm 1 by replacing Steps 2 and 3 with Steps 2’ and 3’ coincides with the original Algorithm 1 and thus becomes exact if and only if the model is a decomposable graphical model.
@@
The fact that Step 2 of Algorithm 1 is immediate by the rational MLE for decomposable graphical models is not surprising, given that we can sample directly by combining urn schemes (see, e.g., [20], Section 9.6). However, an implementation of Algorithm 1 may be easier.
Combining Remark 2.1 and Proposition 2.3, we arrive at the following consequence: if sufficient statistics lie in the boundary of the polyhedral cone generated by the column vectors of $A$, the MLE coincides with the UMVUE and vanishes exactly without bias. In practical terms, we obtain
@@
We summarize the time computational complexity of the algorithms proposed in this paper. We only consider the time complexity because the space complexity, i.e., the memory requirement of the implementation, is small enough that they are of no concern.
@@
$q(\beta;x)={}^{t}(Z_{A}(\beta;x),\partial^{u(1)}Z_{A}(\beta;x),\ldots,\partial^{u(r-1)}Z_{A}(\beta;x)),$
where $\{1,\partial^{u(1)},\ldots,\partial^{u(r-1)}\}$ is the set of the standard monomials of the holonomic ideal of rank $r$, where $\partial^{u(\cdot)}$ is a monomial in $\partial_{j}$, $j\in[m]$. The standard monomials provide the basis of the solutions of the $A$-hypergeometric system. The system of equations (17) can be obtained via a normal form with respect to a Gröbner basis of the holonomic ideal (see, e.g., [18], Theorem 1.4.22). Then, with the contiguity relation (11), we obtain a recurrence relation for the vectors $q(\beta;x)$:
@@ 3 Numerical experiments
We applied the direct sampling algorithm and the approximate algorithm introduced in Section 2.4 to three log-affine models. The first two models are two-way contingency tables. Since the independence model introduced in Example 2.8 is a decomposable graphical model, the UMVUE coincides with the rational MLE by Theorem 2.9, and the original Algorithm 1 works efficiently. The other is a non-independence model, which is not log-linear. In Section 4.2 of [16], the authors applied a previous implementation of the direct sampling algorithm (Algorithm 2.2 of [16]) to the non-independence model. It is exact, but they found it was inefficient due to the computational burden. The last model is the no-three-way interaction model for the three-way contingency table. It is log-linear, but not a decomposable graphical model, so the UMVUE and the MLE never coincide. In Section 4.3 of [16], the authors discussed the model but did not experiment since it involves the recurrence relation (18) for a vector of 81 dimensions, and its computation was prohibitive. Because of the above reasons, we applied the approximate direct sampling algorithm to the last two models.
@@
To evaluate the performance of the Metropolis algorithm, it is first necessary to assess the convergence of the Metropolis chain to the stationary distribution. It is a nontrivial task with various arguments (see, e.g., [19], Section 12.2). In this study, taking advantage of the situation where we have another feasible sampling algorithm, we assessed the convergence of the Metropolis chain using the direct sampling algorithm.
@@
Here, we generated many contingency tables for $\hat{p}_{\rm D}$ in advance so that the empirical distribution of the direct sampling algorithm gives the best possible approximation of the truth (19). If we use the (exact) direct sampling algorithm, (20) approximates the total variation distance between the distribution of chi-square values taken from the Metropolis chain and that of the stationary distribution. If otherwise, we can still assess the convergence of the Metropolis chain because the total variation distance should converge to a non-zero value.
@@
We conducted all computations in the R statistical computing environment on floating-point arithmetic. The following timing results were taken on a single core of the CPU (Intel Core i5-4308U CPU, 2.8GHz) of a machine with 8GB memory.
@@ 3.1 Two-way contingency tables
We investigated the performance of the direct sampling algorithm by exemplifying the following two-way contingency tables with fixed marginal sums for $s=1,2,5$ and $10$:
@@
$u={}^{t}(u_{11},\ldots,u_{15},u_{21},\ldots,u_{25},\ldots,u_{41},\ldots,u_{45}),$
@@
$b={}^{t}(u_{1\cdot},\ldots,u_{4\cdot},u_{\cdot 1},\ldots,u_{\cdot 5})={}^{t}(5s,\ldots,5s,4s,\ldots,4s),$
@@
These tables are unrealistically symmetric, but advantages of this symmetry for the following examinations are that we know the expected counts of each cell for the independent model is uniform, and we expect that the Metropolis chain starting from this tables will mix rapidly. The rational MLE of the expected count is (14) with $i\in[4]$ and $j\in[5]$. For the non-independence model, we set the odds ratios ($x$ in Definition 1.1) as
@@
Then, we assessed the convergence of the Metropolis chain to the stationary distribution. For the Metropolis chain of four pairs of the burn-in and length: $(0,10^{3})$, $(10^{3},10^{4})$, $(10^{4},10^{4})$, and $(10^{4},10^{5})$, respectively, we computed the empirical distribution $\hat{p}_{\rm M}$ of the chi-square values and calculated the total variation distance (20). Here, we set the expected value of each cell in the chi-square to $s$ for both the independence and non-independence models with the odds ratios (21). We show the averages of the results obtained by repeating this procedure 10 times in Table 1. The pair $(0,10^{3})$ gave the poorest results because only a few contingency tables appeared in a Metropolis chain, and they cannot approximate the stationary distribution. The pairs $(10^{3},10^{4})$ and $(10^{4},10^{4})$ were equally good, which implies that the Metropolis chain converged after 10,000 steps. For the pair $(10^{4},10^{5})$, the approximation is even better because the empirical distribution approximates the stationary distribution better by many contingency tables because of the Glivenko–Cantelli theorem. These trends were similar for non-independence models. Another important observation is that for approximate sampling of the non-independence models, the total variation distance did not decrease as $s$ increases. While one would expect the total variation distance to decrease as the bias of the MLE decreases asymptotically, the reverse trend suggests that the Monte Carlo error is relatively large compared to the bias and that the impact of the bias is small in typical settings. Indeed, the total variation distances showed a similar trend for the direct sampling from the independence model.
@@
Given the above results, we drew 10,000 tables after the burn-in of 10,000 steps for the Metropolis chain. We computed the ESS of the 10,000 tables drawn by the Metropolis algorithm and drew the same number of tables using the direct sampling algorithm. Table 2 shows the averages of the results obtained by repeating this procedure 100 times. It shows the times it took to draw 10,000 tables using the Metropolis algorithm (including the burn-in), the time it took to draw the tables of the number of the ESS using the direct sampling algorithm, and one table using the direct sampling algorithm. For the independence model, the direct sampling algorithm was uniformly efficient than the Metropolis algorithm. For the non-independence model, although the direct sampling algorithm was inefficient than the Metropolis algorithm, even when the total count is 200 ($s=10$), which is the most computationally expensive for the direct sampling algorithm, the direct sampling algorithm took only about six times as long as the Metropolis algorithm.
@@ 3.2 No-three-way interaction model of three-way table
$b={}^{t}($ $u_{11\cdot},u_{12\cdot},u_{13\cdot},u_{21\cdot},u_{22\cdot},u_{23\cdot},u_{1\cdot 1},u_{1\cdot 2},u_{1\cdot 3},u_{2\cdot 1},u_{2\cdot 2},u_{2\cdot 3},$
$u_{\cdot 11},u_{\cdot 12},u_{\cdot 13},u_{\cdot 21},u_{\cdot 22},u_{\cdot 23},u_{\cdot 31},u_{\cdot 32},u_{\cdot 33})$
=($={}^{t}($ $3s,\ldots,3s,2s,\ldots,2s),$
@@ A Markov basis consists of [6]
$\begin{bmatrix}1i_{2}i_{3}\\ 1j_{2}j_{3}\\ 2i_{2}j_{3}\\ 2j_{2}i_{3}\end{bmatrix}-\begin{bmatrix}1i_{2}j_{3}\\ 1j_{2}i_{3}\\ 2i_{2}i_{3}\\ 2j_{2}j_{3}\end{bmatrix},\quad i_{2}\neq j_{2},~{}i_{3}\neq j_{3}\in[3],$
which represents nine vectors with the components of $u$ in the first bracket are $+1$, those in the seconds brackets are $-1$, and the other components are 0, and six vectors with 12 non-zero elements including
@@
where $(i_{1},i_{2},i_{3})$ is a permutation of $(1,2,3)$ and determines $(j_{1},j_{2},j_{3})$. For example, if $(i_{1},i_{2},i_{3})=(1,3,2)$, then $(j_{1},j_{2},j_{3})=(2,1,3)$.
@@
and $p(z)=0$ if $z\notin\{0,8,12\}$. The closeness to the true distribution (22) of the chi-square values demonstrates that the MLE performs a good approximation of the UMVUE. The pairs $(10^{4},10^{4})$ and $(10^{5},10^{4})$ were equally good, which implies that the Metropolis chain converged after 100,000 steps. An important observation is that the total variation distance of the pair $(10^{5},10^{5})$ decreases as $s$ increase except $s=1$ (the Monte Carlo error was small for $s=1$ since we used not the empirical distribution $\hat{p}_{D}$ but the true distribution (22)). Unlike the case of the two-way contingency tables (Table 1), the bias decrease of the MLE is evident.
@@
Given the above results, we drew 10,000 tables after the burn-in of 100,000 steps for the Metropolis chain. We computed the ESS of the 10,000 tables drawn by the Metropolis algorithm and drew the same number of tables using the direct sampling algorithm. Table 4 shows the averages of the results obtained by repeating this procedure 10 times. It shows the times it took to draw 10,000 tables using the Metropolis algorithm (including the burn-in), the time it took to draw the tables of the number of the ESS using the approximate direct sampling algorithm, and one table using the direct sampling algorithm. The approximate direct sampling algorithm was uniformly inefficient than the Metropolis algorithm. When the total count is 180 ($s=10$), it took about 36 times as long as the Metropolis algorithm.
Table 4: The times to draw 10,000 tables by the Metropolis algorithm, the effective sample sizes (ESS) of the 10,000 tables, the times to draw the same number of tables by the direct sampling algorithm, and the times per single table. Times are in seconds.
$s$ ESS Metropolis Direct 1 table

### v3

@@ 2.4 The approximate algorithm without using connection matrices
In this subsection, we propose to use Algorithm 1 with Steps 2 and 3 replaced by Steps 2’ and 3’, respectively. We call the resulting algorithm the approximate direct sampling algorithm. From Lemma 2.4, when the MLE is not the UMVUE, i.e., it is not the rational MLE, then there exists a bias that disappears asymptotically. Thus the resulting algorithm is approximate in the sense that it is not exact.
@@
For an undirected discrete graphical model, the approximate algorithm obtained from Algorithm 1 by replacing Steps 2 and 3 with Steps 2’ and 3’ coincides with the original Algorithm 1 if and only if the model is a decomposable graphical model. Then, the algorithm is exact.
@@
The fact that Step 2 of Algorithm 1 is immediate by the rational MLE for decomposable graphical models is not surprising, given that we can sample directly by combining urn schemes (see, e.g., [22, Section 9.6]). However, an implementation of Algorithm 1 may be easier.
Combining Remark 2.1 and Proposition 2.5, we arrive at the following consequence: if sufficient statistics lie in the boundary of the polyhedral cone generated by the column vectors of $A$, the MLE coincides with the UMVUE and vanishes exactly without bias. In practical terms, we obtain the following proposition.
@@
Proof.
(Markov property) Even in the algorithm where Steps 2 and 3 are replaced by Steps 2’ and 3’, the transition probability remains determined by the current state, so the Markov property holds.
(agreement of positivity of transition probabilities) Proposition 2.5 asserts that when a component of the UMVUE vanishes, that component of the MLE also vanishes, and vice versa. Therefore, for a transition in which the transition probability of Algorithm 1 is 0, the transition probability of the approximate algorithm is also 0, and vice versa. Furthermore, for a transition in which the transition probability of Algorithm 1 is positive, the transition probability of the approximate is also positive, and vice versa.
(agreement of supports of sample paths) Let $(j_{1},...,j_{n})$ denote a sample path obtained with positive probability in Algorithm 1, and $p_{j_{t},j_{t+1}}$, t=1,2,,n1t=1,2,...,n-1 denote the transition probabilities of each step. Since the algorithm produces a Markov chain, the probability of obtaining the sample path is given by the product of these, so $p_{j_{t},j_{t+1}}$ must all be positive. The transition probability of each step of the approximate algorithm is expressed as $p^{\prime}_{j_{t},j_{t+1}}$, t=1,2,,n1t=1,2,...,n-1. In this case, from the agreement of positivity of transition probabilities shown above, we can see that the probability of obtaining the sample path $(j_{1},...,j_{n})$, namely, the product of $p^{\prime}_{j_{t},j_{t+1}}$, t=1,2,,n1t=1,2,...,n-1, is positive. Conversely, a sample path in the approximate algorithm is also obtained with a positive probability in Algorithm 1. Thus, we conclude that whether a certain sample path is positive or zero agrees between Algorithm 1 and the approximate algorithm. Since the Markov lattice is the support of sample paths of Algorithm 1, it is also the support of sample paths of the approximate algorithm. ∎
Remark 2.16.
In general, the MLE has asymptotic normality; that is, $\sqrt{n}(\hat{\mu}-\mu)$ converges in distribution to the normal distribution of expectation zero, where $n$ is the total number of counts. As a result, the MLE is asymptotically unbiased. For the covariance of the normal distribution, see [12], page 78. However, this does not mean that the approximate algorithms is unbiased asymptotically when $n$ is large. This is because the algorithm proceeds by reducing the total number of counts in the contingency table with each transition (see the tables in the first line of (28)), so the total number of counts in the contingency tables at later steps of the algorithm becomes small. See Section 3 for numerical examples.
@@
We summarize the computational cost of the algorithms proposed in this paper. We only consider the time complexity because the space complexity, i.e., the memory requirement of the implementation, is small enough that they are of no concern.
@@
$q(\beta;x)=(Z_{A}(\beta;x),\partial^{u(1)}Z_{A}(\beta;x),\ldots,\partial^{u(r-1)}Z_{A}(\beta;x))^{\top},$
where $\{1,\partial^{u(1)},\ldots,\partial^{u(r-1)}\}$ is the set of the standard monomials of the holonomic ideal of rank $r$, where $\partial^{u(\cdot)}$ is a monomial in $\partial_{j}$, $j\in[m]$. The standard monomials provide the basis of the solutions of the $A$-hypergeometric system. The system of equations (41) can be obtained via a normal form with respect to a Gröbner basis of the holonomic ideal (see, e.g., [21, Theorem 1.4.22]). Then, with the contiguity relation (33), we obtain a recurrence relation for the vectors $q(\beta;x)$:
@@ 3 Numerical experiments
We applied the direct sampling algorithm and the approximate algorithm introduced in Section 2.4 to three log-affine models. The first two models are two-way contingency tables. Since the independence model introduced in Example 1.3 is a decomposable graphical model, the UMVUE coincides with the rational MLE by Theorem 2.9, and the original Algorithm 1 works efficiently. The other is a non-independence model, which is not log-linear. In Section 4.2 of [17], the authors applied a previous implementation of the direct sampling algorithm (Algorithm 2.2 of [17]) to the non-independence model. It is exact, but they found it was inefficient due to the computational load. The last model is a no-three-way interaction model for the three-way contingency tables. It is log-linear, but not a graphical model (the edges corresponding to the two-way interactions make the complete graph of the three vertices, but the model has no three-way interaction), and one can see that the UMVUE and the MLE do not coincide. In Section 4.3 of [17], the authors discussed the model but did not experiment since it involves the recurrence relation (42) for a vector of 81 dimensions, and its computation was prohibitive.
@@
To evaluate the performance of the Metropolis algorithm, it is first necessary to assess the convergence of the Metropolis chain to the stationary distribution. It is a nontrivial task with various arguments (see, e.g., [20, Section 12.2]). In this study, taking advantage of the situation where we have another feasible sampling algorithm, we assessed the convergence of the Metropolis chain using the direct sampling algorithm.
@@
Here, we generated many contingency tables for $\hat{p}_{\rm D}$ in advance so that the empirical distribution of the direct sampling algorithm gives the best possible approximation of the truth (43). If we use the (exact) direct sampling algorithm, (44) approximates the total variation distance between the distribution of chi-square values taken from the Metropolis chain and that of the stationary distribution. Otherwise, we can still assess the convergence of the Metropolis chain because the total variation distance should converge to a non-zero value.
@@
The results of the numerical experiments described in this section reveal the following:
  • (1)
    For the non-independence model of two-way contingency tables, which is a decomposable graphical model, the direct sampling algorithm runs faster than the Metropolis algorithm.
  • (2)
    For general log-affine models, the direct sampling algorithm requires the calculation of transition probabilities. It is different from the Metropolis algorithm, where we can use the same Markov basis as the corresponding log-linear model ($x=1$ in Definition 1.1). For a non-independence model of two-way contingency tables, the approximate direct sampling algorithm is several tens of times faster than the exact direct sampling algorithm. Although it needs a larger amount of time than the Metropolis algorithm in terms of the effective sample size, it can be calculated in a similar amount of time. If the number of contingency tables is fixed, as is often done when using the Metropolis algorithm, the approximate algorithm achieves sampling closer to the true distribution than the Metropolis algorithm.
  • (3)
    As an example of a non-graphical log-linear model, we examined the no-three-way interaction model of three-way contingency tables. Its Markov basis is known. The exact direct sampling algorithm was infeasible because of the computational burden. The approximate direct sampling algorithm required several tens of times longer than the Metropolis algorithm in terms of the effective sample size. However, note that the approximate direct sampling algorithm works whenever the MLE is available. The Metropolis algorithm cannot be applied unless we know a Markov basis.
We conducted all computations in the R statistical computing environment, using floating-point arithmetic, except for the results shown at the end of Section 3.1. The following timing results were taken on a single core of the CPU (Intel Core i5-4308U CPU, 2.8GHz) of a machine with 8GB memory.
@@ 3.1 Two-way contingency tables
We investigated the performance of the direct sampling algorithm by exemplifying the two-way contingency tables. The independence model is a log-linear decomposable model (see Example 2.11), while the non-independence model is a log-affine model but not a log-linear model. We investigated following two-way contingency tables with fixed marginal sums for $s=1,2,5$ and $10$:
@@
$u=(u_{11},\ldots,u_{15},u_{21},\ldots,u_{25},\ldots,u_{41},\ldots,u_{45})^{\top},$
@@
$b=(u_{1\cdot},\ldots,u_{4\cdot},u_{\cdot 1},\ldots,u_{\cdot 5})^{\top}=(5s,\ldots,5s,4s,\ldots,4s)^{\top},$
@@
These tables are unrealistically symmetric, but advantages of this symmetry for the following investigations are that we know the expected counts of each cell for the independent model is uniform, and we expect that the Metropolis chain starting from this tables will mix rapidly. The rational MLE of the expected count is (37) with $i\in[4]$ and $j\in[5]$. For the non-independence model, we set the odds ratios ($x$ in Definition 1.1) as
@@
Then, we assessed the convergence of the Metropolis chain to the stationary distribution. For the Metropolis chain of four pairs of the burn-in and length: $(0,10^{3})$, $(10^{3},10^{4})$, $(10^{4},10^{4})$, and $(10^{4},10^{5})$, respectively, we computed the empirical distribution $\hat{p}_{\rm M}$ of the chi-square values and calculated the total variation distance (44). Here, we set the expected value of each cell in the chi-square to $s$ for both the independence and non-independence models with the odds ratios (45). We show the averages of the results obtained by repeating this procedure 10 times in Table 1. The pair $(0,10^{3})$ gave the poorest results because only a few contingency tables appeared in a Metropolis chain, and they cannot approximate the stationary distribution. The pairs $(10^{3},10^{4})$ and $(10^{4},10^{4})$ were equally good, which implies that the Metropolis chain converged after 10,000 steps. For the pair $(10^{4},10^{5})$, the approximation is even better because the empirical distribution approximates the stationary distribution better by many contingency tables because of the Glivenko–Cantelli theorem. These trends were similar for non-independence models. Another important observation is that for approximate sampling of the non-independence models, the total variation distance did not decrease as $s$ increases. Although one might expect the total variation distance to decrease as the bias of the MLE decreases with increasing sample size (see Remark 2.16), the observed result is the opposite: the total variation distance increases. This contrast indicates that the Monte Carlo error dominates the bias, making the impact of the bias comparatively minor. Indeed, the total variation distances showed a similar trend for the direct sampling from the independence model, where the direct sampling is exact without bias.
@@
For the Metropolis algorithm, the same Markov basis can be applied whether the model is independent or not, and the computational cost does not change significantly. On the other hand, for the direct sampling algorithm, the cost of calculating the transition probabilities is an issue. For the independence model, the transition probabilities are given by the rational MLE and can be calculated immediately, but for the non-independence model, calculating the transition probabilities takes time. The purpose here is to evaluate the computational cost.
Given the above results, we drew 10,000 tables after the burn-in of 10,000 steps for the Metropolis chain. We computed the ESS of the 10,000 tables drawn by the Metropolis algorithm and drew the same number of tables using the direct sampling algorithm. Table 2 shows the averages of the results obtained by repeating this procedure 100 times. It shows the times it took to draw 10,000 tables using the Metropolis algorithm (including the burn-in), the time it took to draw the tables of the number of the ESS using the direct sampling algorithm, and one table using the direct sampling algorithm. For the independence model, the direct sampling algorithm was uniformly more efficient than the Metropolis algorithm. For the non-independence model, although the direct sampling algorithm was less efficient than the Metropolis algorithm, even when the total count is 200 ($s=10$), which is the most computationally expensive for the direct sampling algorithm, the direct sampling algorithm took only about six times as long as the Metropolis algorithm.
@@
At the end of this subsection, we compare the accuracy and computational efficiency of the (exact) direct sampling algorithm and the approximate algorithm for the non-independent model. We performed calculations using rational arithmetic in the Risa/Asir computational algebra system [18], since the direct sampling algorithm is implemented on Risa/Asir. Since Algorithm 1 was infeasible for the contingency tables discussed above, we investigated the following two-way contingency tables with fixed marginal sums for $s=1,2,5,10$:
$\begin{array}[]{cccc|c}u_{11}&u_{12}&u_{13}&u_{14}&4s\\ u_{21}&u_{22}&u_{23}&u_{24}&4s\\ u_{31}&u_{32}&u_{33}&u_{34}&4s\\ \hline\cr 3s&3s&3s&3s&12s\end{array}$
with setting the odds ratios as
$\left(\begin{array}[]{cccc}3&2&100/101&1\\ 200/104&200/103&100/102&1\\ 1&1&1&1\\ \end{array}\right).$
Table 3: The times to draw 1,000 and 10,000 tables by the exact direct sampling algorithm and the approximate direct sampling algorithm, respectively, for the non-independence model. The total variation distances between the empirical distributions of chi-square values generated by these algorithms are also shown. The total variation distances between the exact direct sampling algorithm and the Metropolis algorithm are also shown. Times are in seconds.
Exact Approximate Metropolis
$s$ Time Time Distance Distance
1 1581 308 0.019 0.032
2 3496 721 0.051 0.054
5 8513 2171 0.096 0.099
10 18270 5170 0.138 0.151
We generated 1,000 and 10,000 tables using the exact direct sampling algorithm and the approximate direct sampling algorithm, respectively. Table 3 shows the times they took to draw tables, and the results for the approximate direct sampling algorithm are the averages of the results obtained by repeating this procedure 10 times. The approximate direct sampling algorithm performed around 35-50 times faster than the exact direct sampling algorithm. The total variation distances between the empirical distributions of chi-square values generated by the exact and approximate direct sampling algorithms, as well as between the exact direct sampling algorithm and the Metropolis algorithm, are also shown.
For the Metropolis chain, 10,000 tables were drawn following a 10,000-step burn-in. The shown values are averages across 10 repetitions. Although the Metropolis algorithm approximates the stationary distribution well, as shown by the results in Table 2 for larger contingency tables, the approximate direct sampling algorithm gave a distribution closer to that of the exact direct sampling algorithm than the Metropolis algorithm did. It is notable that the approximate algorithm achieves higher accuracy than the Metropolis algorithm for a fixed number of tables, despite the bias introduced by using the MLE to calculate the transition probabilities. This is not surprising, because the effective sample size of tables drawn by the Metropolis algorithm is far smaller than the number of tables. See Table 2 for an example.
@@ 3.2 No-three-way interaction model of three-way table
$b=($ $u_{11\cdot},u_{12\cdot},u_{13\cdot},u_{21\cdot},u_{22\cdot},u_{23\cdot},u_{1\cdot 1},u_{1\cdot 2},u_{1\cdot 3},u_{2\cdot 1},u_{2\cdot 2},u_{2\cdot 3},$
$u_{\cdot 11},u_{\cdot 12},u_{\cdot 13},u_{\cdot 21},u_{\cdot 22},u_{\cdot 23},u_{\cdot 31},u_{\cdot 32},u_{\cdot 33})^{\top}$
$=($ $3s,\ldots,3s,2s,\ldots,2s)^{\top},$
@@ A Markov basis consists of [7]
$\begin{bmatrix}1i_{2}i_{3}\\ 1j_{2}j_{3}\\ 2i_{2}j_{3}\\ 2j_{2}i_{3}\end{bmatrix}-\begin{bmatrix}1i_{2}j_{3}\\ 1j_{2}i_{3}\\ 2i_{2}i_{3}\\ 2j_{2}j_{3}\end{bmatrix},\quad\{(i_{2},j_{2},i_{3},j_{3})\in[3]^{4}:i_{2}<j_{2},i_{3}<j_{3}\},$
which represents nine vectors with the components of $u$ in the first bracket are +1+1, those in the seconds brackets are 1-1, and the other components are 0, and six vectors with 12 non-zero elements including
@@
where $(i_{1},i_{2},i_{3})$ is a permutation of $(1,2,3)$ and determines $(j_{1},j_{2},j_{3})$ by the condition $u\in{\rm Ker}A\cap\mathbb{Z}^{18}$. For example, if $(i_{1},i_{2},i_{3})=(1,3,2)$, then $(j_{1},j_{2},j_{3})=(2,1,3)$.
@@
and $p(z)=0$ if $z\notin\{0,8,12\}$. The closeness to the true distribution (46) of the chi-square values demonstrates that the MLE performs a good approximation of the UMVUE. The pairs $(10^{4},10^{4})$ and $(10^{5},10^{4})$ were equally good, which implies that the Metropolis chain converged after 100,000 steps. An important observation is that the total variation distance of the pair $(10^{5},10^{5})$ decreases as $s$ increase except $s=1$ (the Monte Carlo error was small for $s=1$ since we used not the empirical distribution $\hat{p}_{D}$ but the true distribution (46)). This is plausible, since as increasing the sample size reduces the bias of the MLE (see Remark 2.16). Unlike the case of the two-way contingency tables (Table 1), the bias decrease is evident.
@@
Given the above results, we drew 10,000 tables after the burn-in of 100,000 steps for the Metropolis chain. We computed the ESS of the 10,000 tables drawn by the Metropolis algorithm and drew the same number of tables using the direct sampling algorithm. Table 5 shows the averages of the results obtained by repeating this procedure 10 times. It shows the times it took to draw 10,000 tables using the Metropolis algorithm (including the burn-in), the time it took to draw the tables of the number of the ESS using the approximate direct sampling algorithm, and one table using the direct sampling algorithm. The approximate direct sampling algorithm was uniformly less efficient than the Metropolis algorithm. When the total count is 180 ($s=10$), it took about 36 times as long as the Metropolis algorithm.
Table 5: The times to draw 10,000 tables by the Metropolis algorithm, the effective sample sizes (ESS) of the 10,000 tables, the times to draw the same number of tables by the approximate direct sampling algorithm, and the times per single table. Times are in seconds.
$s$ ESS Metropolis Approx.-Direct 1 table